OpenStudy (anonymous):

need to get points to graph by solving equation y-5x=0

OpenStudy (anonymous):

y-5x=0 y=5x <- Canonical form/intercept form Choose any x you want (0 and 1 are good candidate), plot for y.

OpenStudy (anonymous):

u can plot it on geogebra software and then post it here we would check if u r having problem with graphing

OpenStudy (anonymous):

it is regular function where y depends on x. It is just given in another form. therefore after you will set it normally, you will get y=5x. Now you have system of coordinates with y-axes and x-axes. for each value of x you have your y. For example. when x=0, y=5*0=0. So the point will be (0:0). when x=1, y=5*1=5, so you got point (1:5) etc.